Use the given zero to find the remaining zeros of the function. f(x)=x^3-9x^2+36x-324; zero 6i
Answer by stanbon(75887)

Use the given zero to find the remaining zeros of the function. f(x)=x^3-9x^2+36x-324; zero 6i
----
Since the coefficients are all REAL Numbers, -6i is also a zero.
--------
(x-6i)(x+6i) = x^2+36 must be a factor of f(x).
---
3rd factor is (x-9)
3rd zero is x = 9
=======================
Cheers,
Stan H.
----